# SATA: Safe and Adaptive Torque-Based Locomotion Policies Inspired by Animal Learning

### v1

@@ Abstract
Despite recent advances in learning-based controllers for legged robots, deployments in human-centric environments remain limited by safety concerns. Most of these approaches use position-based control, where policies output target joint angles that must be processed by a low-level controller (e.g., PD or impedance controllers) to compute joint torques. Although impressive results have been achieved in controlled real-world scenarios, these methods often struggle with compliance and adaptability when encountering environments or disturbances unseen during training, potentially resulting in extreme or unsafe behaviors. Inspired by how animals achieve smooth and adaptive movements by controlling muscle extension and contraction, torque-based policies offer a promising alternative by enabling precise and direct control of the actuators in torque space. In principle, this approach facilitates more effective interactions with the environment, resulting in safer and more adaptable behaviors. However, challenges such as a highly nonlinear state space and inefficient exploration during training have hindered their broader adoption. To address these limitations, we propose SATA, a bio-inspired framework that mimics key biomechanical principles and adaptive learning mechanisms observed in animal locomotion. Our approach effectively addresses the inherent challenges of learning torque-based policies by significantly improving early-stage exploration, leading to high-performance final policies. Remarkably, our method achieves zero-shot sim-to-real transfer, eliminating the need for additional fine-tuning on hardware. Our experimental results indicate that SATA demonstrates remarkable compliance and safety, even in challenging environments such as soft/slippery terrain or narrow passages, and under significant external disturbances (e.g., pushing/pulling/pressing on the robot, or manually moving individual legs). These results highlight its potential for practical deployments in human-centric and safety-critical scenarios.
@@ I Introduction
Most existing RL-based methods rely on position control [6, 7]. In this approach, neural networks output target joint positions, which are subsequently translated into joint torque commands through a low-level proportional–derivative (PD) controller. Position-based policies are simple and easy to train, as they abstract away the complexities of actuation physics and dynamics. However, this simplicity limits the policy’s capacity to explore fine-grained and dynamic behaviors, thereby reducing its adaptability and generalization to unseen challenges in real-world environments. For instance, position-based policies trained on rigid terrains in simulation often struggle to generalize to deformable environments in the real world due to their out-of-distribution nature. As a result, these methods usually rely heavily on terrain randomization, where the agent trains in large amounts of potential environments to maximize the chances that the final policy may handle real-world environments [8, 9]. Furthermore, position-based strategies exhibit notable limitations in compliance. Lacking the ability to directly regulate joint torques, these methods often lead to overreactions to external disturbances. For example, if one of the robot’s legs suddenly gets stuck, a position-based policy may aggressively attempt to force the actuator into the commanded position. This can cause the motors to generate excessively high torques, potentially destabilizing or even damaging the robot, or creating safety hazards to its surroundings, such as nearby objects or even humans.
@@ II Related Work
Figure 2: Overview of our SATA Framework. Dotted lines indicate parts used only during training, while solid lines those used during both training and deployment. Our framework is mainly composed of 1) a Biomechanical Model (Orange) to ensure the generation of smooth, practical actuator commands $\tau$ while informing the policy of the current actuator state, and 2) a Growth Model (Green) to help the neural network train a more robust and generalizable policy by gradually adapting rewards $r_{\textit{growth}}$, control frequency $f_{\textit{policy}}$, and torque limits $\tau_{\textit{limit}}$ during training. Finally, we train a state estimator for real world deployment using simulated IMU data and temporal proprioception observations (Grey), to help condition our policy on the (estimated) current robot velocity.
Quadruped controllers have greatly benefited from the development of deep reinforcement learning (DRL) in recent years, allowing agents to learn impressive controllers [6, 15, 16, 17] that would otherwise require the design and solving of non-linear optimization problems, which often involves approximations that cannot be neglected in real-world settings [18, 19, 20, 21]. However, the problem of sampling efficiency still hinders its application. Early approaches employed by  Peng et al. [22] combined imitation learning with DRL to solve this problem, demonstrating gaits similar to the training set while adapting to different terrain or disturbances. This approach allows for the training of adaptive quadruped controllers in highly dynamic acrobatic tasks, but its performance is limited by the dataset quality and robot deployment remains nontrivial. Addressing these problems, Hwangbo et al.  [23] proposed a method to train deployable policies in simulation by introducing an actuator network and domain randomization. Their learned policies can yield different gaits or recover from fallen positions. To help the robot travel in challenging terrains,  Lee et al. [4] introduced a teacher-student framework for quadrupedal robots, allowing them to traverse complex terrains without any visual feedback. To aid locomotion with environmental information, [5] built upon [4] by integrating additional sensors and employing an attention-based recurrent encoder to fuse proprioceptive and exteroceptive inputs. This resulted in a robust and fast legged motion controller for navigating challenging terrains. While exteroceptive inputs can enable more informed decision-making, a highly compliant and robust base policy—capable of operating effectively without vision—remains crucial for reliable real-world deployment. Moreover, reliance on exteroception introduces additional challenges, such as the sim-to-real gap, where sensor noise, latency, and real-world variations degrade performance.
@@ III-A1 Activation Model
Output by our policy network, the action signal $a_{s}$ first passes through the activation model. This model functions similarly to motor neurons [55, 56], transforming the action signal into the corresponding activation signal, $\alpha_{\text{current}}$:
@@
In this formulation, $\gamma$ serves as a weighting factor. This ensures a smooth temporal evolution of activation signals, which promotes continuity and smoothness in movements.
@@ III-A3 Internal State Model
Our internal state model does not directly participate in torque calculation but serves as a feedback mechanism to provide the robot with more information. In this module, we simulate the fatigue mechanism observed in animals to construct a dynamic state indicator that quantifies the accumulation and recovery of fatigue during the robot’s operation. The fatigue indicator is directly influenced by the joint torque intensity and evolves dynamically over time, following the equation:
@@ III-B1 Observation and Action Space for Actor
Other components of $o_{t}$ include $w_{t}\in\mathbb{R}^{3}$, the angular velocity of the robot base, and $g_{t}\in\mathbb{R}^{3}$, the gravity direction vector in the body frame. These quantities aid in orientation estimation and maintaining balance. Additionally, $v_{\text{cmd}}\in\mathbb{R}^{3}$ specifies the desired linear and angular velocities, while $q\in\mathbb{R}^{12}$, $\dot{q}\in\mathbb{R}^{12}$, $\tau\in\mathbb{R}^{12}$, and $\zeta\in\mathbb{R}^{12}$ represent the joint states, joint torques, and fatigue state, as described earlier.
@@ III-B2 Reward Design
TABLE I: Reward components and their corresponding weights ($dt=0.005$)
@@ IV Growth-Based Training
Due to the highly nonlinear characteristics of the torque space, training a torque-based policy poses greater challenges compared to a position-based policy, especially during early exploration. To address this, we introduce a growth mechanism that progressively unlocks the robot’s capabilities, dynamically adjusts training objectives, and incrementally increases the control frequency, facilitating more efficient policy training and improving the generalization ability of the learned policy.
@@ IV-A Implementation of the Growth Mechanism
Instead of granting the policy full access to the action space from the start of training, we propose that partially limiting the robot’s abilities can promote more efficient exploration. Additionally, gradually increasing the control frequency simplifies exploration and mitigates the problem of delayed reward during the early stages of training. To unify these components, we introduce a time-dependent general scale $G(t)$, derived from the Gompertz model, a well-established framework for modeling growth:
@@ IV-A1 Torque Limits and Control Frequency Adjustment
Using $G(t)$, we dynamically update the torque limits ($\tau_{\text{limit}}$) and control frequency ($f_{policy}$) during training. These updates enable the robot to gradually get access to its full operational capabilities:
@@ IV-B Training Details
During the training process, the control frequency and torque limit are progressively increased, asymptotically approaching their full capacity but never fully reaching it, as described in Eqs. 6 and 7. Specifically, we set $k=0.00003$, $t_{0}=24000$, $\tau_{\text{start}}=7.05\,\mathrm{Nm}$, $\tau_{\text{end}}=23.5\,\mathrm{Nm}$, $f_{\text{start}}=100\,\mathrm{Hz}$, and $f_{\text{end}}=200\,\mathrm{Hz}$. These parameters enable the robot to achieve approximately 50% of its full capability after around 1.5k training episodes and approximately 70% after about 2.5k episodes.
For deployment, the robot’s full capacity is released by setting both the control frequency and torque limit to their maximum values, $f_{\text{end}}$ and $\tau_{\text{end}}$, respectively.
@@ V-A1 Ablation Study
As shown in Fig. 3, the biomechanical model plays a critical role in enabling natural and stable locomotion. When this biomechanical model is removed, the robot converges to unnatural gaits, such as three-legged support patterns, which reduce stability and limit energy efficiency. This highlights the importance of the biomechanical model and feedback mechanisms in smoothing motion commands and preventing suboptimal convergence. Meanwhile, the absence of the growth mechanism leads to limited generalization to higher velocity commands. While SATA w/o Growth demonstrates stable and smooth locomotion at low speeds ($v_{x}=0.7$), it fails to track higher velocity commands stably $v_{x}=1.5$, the maximum velocity encountered during training. In contrast, the SATA framework performs robustly across all tested velocity commands. Not only does it maintain stability over the full range of velocity commands seen during training, it also generalizes well to tracking $v_{x}=2.0$, a velocity well over this range. This indicates that the inclusion of progressive training adjustments improves the policy’s ability to handle higher velocities, demonstrating SATA’s superior generalizability, and allowing it to become more robust to out-of-distribution scenarios. In particular, these results suggest that the complementary roles of the biomechanical model help ensure stable and natural motion, and the growth mechanism enhance the policy’s adaptability and robustness under diverse conditions.
@@ V-A2 Robustness to Single-Leg Failure
TABLE III: Performance Comparison between Our Method and Baselines Across Different Robustness Tests (5 Trials per Test)
@@ V-B Lab Level Experiments
To validate the effectiveness of our approach, we deployed it on a Unitree Go2 quadruped robot in real-world scenarios. We also compared its performance against several baseline methods, including Unitree’s built-in, MPC-based controller, a vanilla learning position-based policy, and a well-known learning position-based policy, WalkTheseWay [9]. Experimental results demonstrated that our approach exhibited robust performance across various scenarios, including handling unexpected disturbances and navigating unseen environments not encountered during training. In multiple evaluations, as summarized in Table III, our method consistently outperformed the baseline approaches across all robustness tests (please refer to our associated video for details).
@@ V-B3 Kicking and Stomping
Figure 7: Walking under external downward forces/presses. The blue line is the actual torque command given to the motors before the processing done by our biomechanical model, while the orange dotted line is the raw action output of the policy. The robot can continue progressing forward even when external force press it to the ground. Note that the biomechanical model is crucial in ensuring the safety of the robot especially during these disturbances, limiting excessive torque output and mitigating oscillations.

### v2

@@ Abstract
Despite recent advances in learning-based controllers for legged robots, deployments in human-centric environments remain limited by safety concerns. Most of these approaches use position-based control, where policies output target joint angles that must be processed by a low-level controller (e.g., PD or impedance controllers) to compute joint torques. Although impressive results have been achieved in controlled real-world scenarios, these methods often struggle with compliance and adaptability when encountering environments or disturbances unseen during training, potentially resulting in extreme or unsafe behaviors. Inspired by how animals achieve smooth and adaptive movements by controlling muscle extension and contraction, torque-based policies offer a promising alternative by enabling precise and direct control of the actuators in torque space. In principle, this approach facilitates more effective interactions with the environment, resulting in safer and more adaptable behaviors. However, challenges such as a highly nonlinear state space and inefficient exploration during training have hindered their broader adoption. To address these limitations, we propose Safe and Adaptive Torque-based locomotion policies inspired by Animal learning(SATA), a bio-inspired framework that mimics key biomechanical principles and adaptive learning mechanisms observed in animal locomotion. Our approach effectively addresses the inherent challenges of learning torque-based policies by significantly improving early-stage exploration, leading to high-performance final policies. Remarkably, our method achieves zero-shot sim-to-real transfer, eliminating the need for additional fine-tuning on hardware. Our experimental results indicate that SATA demonstrates remarkable compliance and safety, even in challenging environments such as soft/slippery terrain or narrow passages, and under significant external disturbances (e.g., pushing/pulling/pressing on the robot, or manually moving individual legs). These results highlight its potential for practical deployments in human-centric and safety-critical scenarios.
@@ I Introduction
Most existing RL-based methods rely on position control [6, 7]. In this approach, neural networks output target joint positions, which are subsequently translated into joint torque commands through a low-level proportional-derivative (PD) controller. Position-based policies are simple and easy to train, as they abstract away the complexities of actuation physics and dynamics. However, this simplicity limits the policy’s capacity to explore fine-grained and dynamic behaviors, thereby reducing its adaptability and generalization to unseen challenges in real-world environments. For instance, position-based policies trained on rigid terrains in simulation often struggle to generalize to deformable environments in the real world due to their out-of-distribution nature. As a result, these methods usually rely heavily on terrain randomization, where the agent trains in large amounts of potential environments to maximize the chances that the final policy may handle real-world environments [8, 9]. Furthermore, position-based strategies exhibit notable limitations in compliance. Lacking the ability to directly regulate joint torques, these methods often lead to overreactions to external disturbances. For example, if one of the robot’s legs suddenly gets stuck, a position-based policy may aggressively attempt to force the actuator into the commanded position. This can cause the motors to generate excessively high torques, potentially destabilizing or even damaging the robot, or creating safety hazards to its surroundings, such as nearby objects or even humans.
@@ II Related Work
Figure 2: Overview of our SATA Framework. Dotted lines indicate parts used only during training, while solid lines indicate those used during both training and deployment. Our framework is mainly composed of 1) a Biomechanical Model (Orange) to ensure the generation of smooth, practical actuator commands $\tau$ while informing the policy of the current actuator state, and 2) a Growth Model (Green) to help the neural network train a more robust and generalizable policy by gradually adapting rewards $r_{\textit{growth}}$, control frequency $f_{\textit{policy}}$, and torque limits $\tau_{\textit{limit}}$ during training. Finally, we train a state estimator for real world deployment using simulated IMU data and temporal proprioception observations (Grey), to help condition our policy on the (estimated) current robot velocity.
Quadruped controllers have greatly benefited from the development of deep reinforcement learning (DRL) in recent years, allowing agents to learn impressive controllers [6, 15, 16, 17] that would otherwise require the design and solving of non-linear optimization problems, which often involves approximations that cannot be neglected in real-world settings [18, 19, 20, 21]. However, the problem of sampling efficiency still hinders its application. Early approaches employed by Peng et al. [22] combined imitation learning with DRL to solve this problem, demonstrating gaits similar to the training set while adapting to different terrain or disturbances. This approach allows for the training of adaptive quadruped controllers in highly dynamic acrobatic tasks, but its performance is limited by the dataset quality and robot deployment remains nontrivial. Addressing these problems, Hwangbo et al. [23] proposed a method to train deployable policies in simulation by introducing an actuator network and domain randomization. Their learned policies can yield different gaits or recover from fallen positions. To help the robot travel in challenging terrains, Lee et al. [4] introduced a teacher-student framework for quadrupedal robots, allowing them to traverse complex terrains without any visual feedback. To aid locomotion with environmental information, [5] built upon [4] by integrating additional sensors and employing an attention-based recurrent encoder to fuse proprioceptive and exteroceptive inputs. This resulted in a robust and fast legged motion controller for navigating challenging terrains. While exteroceptive inputs can enable more informed decision-making, a highly compliant and robust base policy - capable of operating effectively without vision - remains crucial for reliable real-world deployment. Moreover, reliance on exteroception introduces additional challenges, such as the sim-to-real gap, where sensor noise, latency, and real-world variations degrade performance.
@@ III-A1 Activation Model
Output by our policy network, the action signal $a_{s}$ first passes through the activation model [55]. This model functions similarly to motor neurons [56, 57], transforming the action signal into the corresponding activation signal, $\alpha_{\text{current}}$:
@@
In this formulation, $\gamma$ serves as a smoothing factor. This ensures a smooth temporal evolution of activation signals, which promotes continuity and smoothness in movements.
@@ III-A3 Internal State Model
Our internal state model does not directly participate in torque calculation but serves as a feedback mechanism to provide the robot with more information. In this module, we simulate the fatigue mechanism observed in animals to construct a dynamic state indicator that quantifies the accumulation and recovery of fatigue during the robot’s operation [62, 63]. The fatigue indicator is directly influenced by the joint torque intensity and evolves dynamically over time, following the equation:
@@ III-B1 Observation and Action Space for Actor
Other components of $o_{t}$ include $w_{t}\in\mathbb{R}^{3}$, the angular velocity of the robot base, and $g_{t}\in\mathbb{R}^{3}$, the gravity direction vector in the body frame. These quantities aid in orientation estimation and maintaining balance. Additionally, $v_{\text{cmd}}\in\mathbb{R}^{3}$ specifies the desired linear and angular velocities, while $q\in\mathbb{R}^{12}$, $\dot{q}\in\mathbb{R}^{12}$, $\tau\in\mathbb{R}^{12}$, and $\zeta\in\mathbb{R}^{12}$ represent the joint states, joint velocities, joint torques, and fatigue state, as described earlier.
@@ III-B2 Reward Design
TABLE I: Reward components and weights ($dt=0.005$)
@@ IV Growth-Based Training
Due to the highly nonlinear nature of the torque space, training a torque-based policy poses greater challenges than a position-based one, especially during early-stage exploration. To address this, we propose a biologically inspired growth mechanism that mimics animal development by progressively unlocking the robot’s physical capabilities, dynamically adapting reward functions, and gradually increasing control frequency. This process facilitates more efficient policy learning while preserving stability and promoting generalization.
While related in spirit to curriculum learning [64, 65, 66, 67] and progressive learning [68, 69] approaches - which typically increase task difficulty over time - our method maintains a fixed task throughout training. Instead of staging increasingly complex goals, we focus on enhancing the agent’s embodiment by gradually expanding what it is physically allowed to do. This leads to deeper exploration and reduces the risk of suboptimal shortcuts, such as exploiting a single powerful joint. Our strategy also aligns conceptually with reward scheduling techniques, where the learning signal evolves in tandem with the agent’s growing capabilities [70, 71]. For instance, as the agent develops from basic contact with the ground to full stepping behavior, the reward emphasis naturally shifts to reflect the current developmental stage.
@@ IV-A Implementation of the Growth Mechanism
Instead of granting the policy full access to the action space from the start of training, we propose that partially limiting the robot’s abilities can promote more efficient exploration. Additionally, gradually increasing the control frequency simplifies exploration and mitigates the problem of delayed reward during the early stages of training. To unify these components, we introduce a time-dependent general scale $G(t)$, derived from the Gompertz model [72], a well-established framework for modeling growth:
@@ IV-A1 Torque Limits and Control Frequency Adjustment
Using $G(t)$, we dynamically update the torque limits ($\tau_{\text{limit}}$) and control frequency ($f_{policy}$), which are closely linked to growth [73, 74, 75], during training. These updates enable the robot to gradually get access to its full operational capabilities:
@@ IV-B Training Details
During training, control frequency and torque limits are progressively increased, asymptotically approaching their maximum values without fully reaching them, as defined in Eqs.6 and7. For deployment, the robot’s full capacity is restored by setting these parameters to their respective maximums, $f_{\text{end}}$ and $\tau_{\text{end}}$. All hyperparameters related to the growth schedule and biomechanical model are summarized in Table III.
TABLE III: Growth schedule and model hyperparameters
Growth Schedule
$k$ 0.00003 $t_{0}$ 24000 $\tau_{\text{start}}$ 7.05 Nm
$\tau_{\text{end}}$ 23.5 Nm $f_{\text{start}}$ 100 Hz $f_{\text{end}}$ 200 Hz
Biomechanical Model
$\kappa_{\text{scale}}$ 5.0 $\gamma$ 0.6 $\beta$ 0.9
@@ V-A1 Ablation Study
As shown in Fig. 3, the biomechanical model plays a critical role in enabling natural and stable locomotion. When this biomechanical model is removed, the robot converges to unnatural gaits, such as three-legged support patterns, which reduce stability and limit energy efficiency. This highlights the importance of the biomechanical model and feedback mechanisms in smoothing motion commands and preventing suboptimal convergence.
On the other hand, the inclusion of the growth mechanism leads to higher early-stage training efficiency and shows better generalization when tracking out-of-distribution (OOD) velocity commands. As demonstrated in Fig. 5a, SATA significantly outperforms SATA w/o growth in early stages of training, demonstrating the impact of this mechanism in early stage exploration. Moreover, when comparing the cumulative reward of both scenarios under OOD velocity commands ($v_{x}=1.8m/s$) as in Fig. 5b, we can see that our method outperforms SATA w/o growth, demonstrating the impact of the growth mechanism on policy generalization. Upon closer inspection, as in Fig. 3, a risky gait emerges as the policy is commanded to the highest command seen in training and in OOD scenarios, leading to unstable tracking of velocity commands.
In particular, these results suggest that the complementary roles of the biomechanical model help ensure stable and natural motion, and the growth mechanism enhance the policy’s adaptability and robustness under diverse conditions.
Figure 5: Comparison of SATA and SATA w/o Growth. Training rewards (a), without G(t) adaptation, and cumulative rewards in simulation test (b), when commanded to run at 1.8 m/s (slightly OOD).
@@ V-A2 Robustness to Single-Leg Failure
TABLE IV: Performance Comparison between Our Method and Baselines Across Different Robustness Tests (5 Trials per Test)
@@
DeCAP (Pure Torque) 0% 0% 0% 60%
DeCAP (Position-assisted) 80% 40% 0% 80%
@@ V-B Lab Level Experiments
To validate the effectiveness of our approach, we deployed it on a Unitree Go2 quadruped robot in real-world scenarios. We also compared its performance against several baseline methods, including Unitree’s built-in, MPC-based controller, a vanilla learning position-based policy, a well-known learning position-based policy, WalkTheseWay [9], as well as a torque-based policy, DeCAP [11]. Experimental results demonstrated that our approach exhibited robust performance across various scenarios, including handling unexpected disturbances and navigating unseen environments not encountered during training. In multiple evaluations, as summarized in Table IV, our method consistently outperformed the baseline approaches across all robustness tests (please refer to our associated video for details).
@@ V-B1 Safety for Human-Robot Interaction
Figure 8: Walking under external downward forces/presses. The blue line is the actual torque command given to the motors after the processing done by our biomechanical model, while the orange dotted line is the raw action output of the policy. The robot can continue progressing forward even when an external force press it to the ground. Note that the biomechanical model is crucial in ensuring the safety of the robot especially during these disturbances, limiting excessive torque output and mitigating oscillations.
@@ VII Conclusion
VIII ACKNOWLEDGMENTS
This work was supported by the Singapore Ministry of Education Academic Research Fund Tier 1, as well as the National Research Foundation, Singapore (NRF), the Maritime and Port Authority of Singapore (MPA), the Singapore Maritime Institute (SMI) under its Maritime Transformation Programme (Project No. SMI-2022-MTP-01), and the Swiss National Science Foundation, under Grant 200021_197237.